OpenStudy (anonymous):

You own 5 pairs of jeans and want to take 2 of them on vacation. In how many ways can you choose 2 pairs of jeans? (1 point) a. 10 ways b. 15 ways c. 4 ways d. 20 ways I think its D???

OpenStudy (anonymous):

I believe it's a (10) 4 + 3 + 2 + 1 = 10

OpenStudy (anonymous):

* 10, not a 10

OpenStudy (anonymous):

it is five choose two written as \(_5C_2\) or \(\binom{5}{2}\) and computed via \[\binom{5}{2}=\frac{5\times 4}{2}=10\]

OpenStudy (anonymous):

Thx @trogdortheburninator & @satellite73